OpenStudy (anonymous):

Which is the side length of a cube with a volume of 27 m3? A. B. C. D.

OpenStudy (anonymous):

http://prntscr.com/9tmnba

OpenStudy (anonymous):

@m_akram

OpenStudy (anonymous):

@Mehek14

OpenStudy (danjs):

\[\huge V = (side)^3 ~~~so~~~~\sqrt[3]{V} = (Side)\]

OpenStudy (anonymous):

Okay Volume of Cube = (Side)^3

OpenStudy (anonymous):

ok so do i do 3 check mark thing again

OpenStudy (anonymous):

So the Side as Dan said \[\sqrt[3]{v}\]

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

is it 3

OpenStudy (danjs):

yes remember, \[\large \sqrt[3]{x^3}=[x^{3}]^{[1/3]} = x^{3/3} = x\]

OpenStudy (anonymous):

ok it it A

OpenStudy (anonymous):

dan

OpenStudy (danjs):

yeah you are solving y = x^3 for x cube root both sides x = cube root of y

OpenStudy (anonymous):

yes so A right

OpenStudy (purple_pink):

hello again

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

dan

OpenStudy (anonymous):

????

OpenStudy (anonymous):

DANJS